12) You have graphed the cost-volume-profit relationships for a company on a break-even chart after being informed of certain assumptions. Explain how the lines on the chart would change if (a) fixed costs increased over the entire range of activity, (b) selling price per unit decreased, and (c) variable costs per unit increased.
Answer #1

LETS TAKE AN EXAMPLE

SALES PER UNIT 10

VARIABLE COST    3

CONTRIBUTION    7

FIXED COST    5

PROFIT 2

NOW THE CONTRIBUTION TO SALES RATIO IS 7/10*100 =70%

BEP = 5/ 0.7 = 7.14

NOW CASE 1 : fixed COST INCREASES SAY FROM 5 TO 6

BEP = 6/0.7 = 8.57 BREAK EVEN SALES WOULD BE INCRRASED FROM 7.14 TO 8.57 WHICH MEANS SALES LINE IN THE GRAPH WOULD INCREASE.

CASE 2 : SELLING PRICE PER UNIT DECREASED:

SALES COME DOWN FROM 10 TO 9

SALES 9

VARIABLE 3

CONTR 6

FIXED COST 5

PROFIT 1

CSRATIO = 6/9*100 =66.66%

BEP = 5/.6666= 7.5

BREAK EVEN SALES HAS GONE UP FROM 7.14 TO 7.5

CASE 3 VARIABLECOST PER UNIT DECREASED

SALES 10

VC    2

CO 8

FC 5

PROFIT 3

CS RATIO 80%

BEPIS 5/0.8= 6.25

PROFIT INCREASES FROM 2 TO 3 AND BEP REDUCED FROM 7.14 TO 6.25